# Common Trends and Long-Run Multipliers in Nonlinear Structural VARs

### v1

@@ 2.3 High-level assumptions
(vi). The requirement that (LABEL:gradient) hold with $\lVert\boldsymbol{\beta}\rVert\leq\overline{b}$ implies $\boldsymbol{\theta}\circ f_{0}^{-1}$ is Lipschitz, with a Lipschitz constant that is bounded by $\overline{b}$; the converse is also true (see A.1 in A). Excepting cases where $f_{0}$ and $\boldsymbol{\theta}$ interact in a very particular way, this condition signals that our concern here is principally with models in which the equilibrium errors are defined by a function $\boldsymbol{\theta}$ for which $\boldsymbol{\theta}(z)=O(\lVert z\rVert)$ as $\lVert z\rVert\rightarrow\infty$, i.e. which exhibits at most linear asymptotic growth. For example, if $f_{0}=I_{p}$ and $\theta(z)=y-m(x)$, for $z=(y^{\mathsf{T}},x^{\mathsf{T}})^{\mathsf{T}}$, then (iii) requires $m$ to be Lipschitz, and thus that $m(x)=O(\lVert x\rVert)$ as $\lVert x\rVert\rightarrow\infty$, for $r\in\{0,\ldots,p-1\}$.
@@
For the purposes of this paper, the sequence of innovations $\{u_{t}\}$ may be generally treated as a given non-random sequence in $\mathbb{R}^{p}$, since our results are either non-asymptotic, or take limits under the assumption that the shocks cease at some finite horizon $\tau$. However, it will occasionally be useful to indicate the implications of our results in cases where the following holds (as is further developed in a companion paper to the present work).
@@ 2.4 Relationship to the linear VAR
3 Common trends
@@
Before providing our extension of the GJRT to the general setting of (LABEL:nlVAR), we first note that the counterpart of $\boldsymbol{\xi}_{t}$ will continue to follow an autoregression of the form (LABEL:eqerr-linear), but now with time-varying coefficients, in particular with $\boldsymbol{\beta}=\boldsymbol{\beta}_{t}$ now depending on the level of $z_{t}$. While such processes cannot be rendered stationary, we still have a well-defined notion of stability for such processes, which is important for ensuring that $\boldsymbol{\xi}_{t}$ remains of strictly smaller stochastic order than the common trends.
@@
$S_{\chi}^{\mathsf{T}}\coloneqq\begin{bmatrix}-\alpha_{\perp}^{\mathsf{T}}&0\\ 0&I_{r}\end{bmatrix}\begin{bmatrix}0_{p\times r}&I_{p}&\cdots&I_{p}\\ I_{r}&0_{r\times p}&\cdots&0_{r\times p}\end{bmatrix}$
@@
by D.1 in D, which agrees precisely with Johansen (1995, Ch. 4).
@@
4 Attractor spaces
@@
where the fact that $\mathscr{M}_{\mu}$ is indeed a $q$-dimensional manifold follows from the second equality, since $\chi:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ is a homeomorphism.
@@
${\cal S}\subset\mathbb{R}^{p}$ is stable if for every $(u,\mathfrak{z})\in\mathbb{R}^{p}\times\mathbb{R}^{kp}$, there exists a $z_{\infty}\in{\cal S}$ such that $z_{t}(u,\mathfrak{z})\rightarrow z_{\infty}\in{\cal S}$. $z_{\infty}\in\mathbb{R}^{kp}$ is a non-trivial attractor if for every $\mathfrak{z}\in\mathbb{R}^{kp}$, there exists a non-empty $\mathscr{U}(z_{\infty};\mathfrak{z})\subset\mathbb{R}^{p}$ such that $z_{t}(u,\mathfrak{z})\rightarrow z_{\infty}$ for all $u\in\mathscr{U}(z_{\infty};\mathfrak{z})$; we term $\mathscr{U}(z_{\infty};\mathfrak{z})$ the domain of attraction for $z_{\infty}$, given $\mathfrak{z}$. ${\cal S}$ is strictly stable if it is stable and contains only non-trivial attractors.
In other words, if $\mathscr{M}_{\mu}$ is strictly stable, then whatever the current state $\vec{z}_{\tau-1}=\mathfrak{z}$ of the model and the given $z_{\infty}\in\mathscr{M}_{\mu}$, there is always a value for the $\tau$-dated shock that would lead $z_{t}$ to converge to $z_{\infty}$; every element in $\mathscr{M}_{\mu}$ may ultimately be ‘reached’ from $\vec{z}_{\tau-1}$. (Note that this is not as trivial as choosing $u_{\tau}=u$ such that $z_{\tau}=z_{\infty}$ immediately, since what is required is that $z_{t}$ converge to a steady state in which $z_{t}\rightarrow z_{\infty}.)$ Since $\{\boldsymbol{\xi}_{t}\}$ is exponentially stable by 3.1, it follows that $\boldsymbol{\xi}_{t}\rightarrow 0$ as $t\rightarrow\infty$. Hence, noting that $\sum_{s=\tau}^{t}u_{s}=u_{\tau}=u$, we have by that result that
@@
Because the r.h.s. depends on $\mathfrak{z}$ only through $\overline{h}(\mathfrak{z})$, and as $u\in\mathbb{R}^{p}$ varies $\alpha_{\perp}^{\mathsf{T}}[\overline{h}(\mathfrak{z})+u]$ ranges freely over $\mathbb{R}^{q}$, we can induce $z_{\infty}(u;\mathfrak{z})$ to take any desired value $\mathscr{M}_{\mu}$. We thus obtain the following.
@@
As a direct consequence of the preceding, we obtain the following characterisation of the linear combinations of the (reduced form) shocks that do not have permanent effects.
@@
The significance of this result may be explained as follows. Suppose that we have the nonlinear SVAR
$f_{0}(z_{t})$ $=c+\sum_{i=1}^{k}f_{i}(z_{t-i})+u_{t}$ $u_{t}$ $=\Upsilon\varepsilon_{t}$
belonging to class ${\cal M}_{r}$, where $f_{0}$ is suitably normalised such that $\Upsilon$ is identified up to premultiplication by an orthogonal matrix, as it is in the linear SVAR. (The reader may find it helpful to consider the simplified setting in which $f_{0}(z)=\Phi_{0}z$ is linear, so that such a normalisation could be achieved simply by imposing $f_{0}=I_{p}$, though this is not necessary for the argument that follows). Suppose that we partition the structural shocks as
@@
where $\varepsilon_{(1),t}$ takes values in $\mathbb{R}^{r}$, and consists of $r$ structural shocks that are regarded as having no permanent effect on $z_{t}$. Then 4.1 implies that
$\Upsilon\begin{bmatrix}I_{r}\\ 0\end{bmatrix}=\operatorname{sp}\alpha\implies\alpha_{\perp}^{\mathsf{T}}\Upsilon\begin{bmatrix}I_{r}\\ 0\end{bmatrix}$
which provides $r$ long-run identifying restrictions on $\Upsilon$, exactly as in a linear SVAR (see Kilian and Lütkepohl, 2017, Sec. 10.2).
5 Long-run multipliers
Having obtained a characterisation of the attractor space $\mathscr{M}_{\mu}$ for $\{z_{t}\}$, we may finally say something about the permanent effect of a shock at time $t=\tau$. The limiting impulse responses or long-run multipliers can be computed by differentiating $z_{\infty}(u;\mathfrak{z})$ with respect to $u$, i.e. by computing the Jacobian
@@ 5 Long-run multipliers
One concern we might have here, in the general nonlinear case, is the apparent dependence of these long-run multipliers on $\vec{z}_{t-1}=\mathfrak{z}\in\mathbb{R}^{kp}$, which potentially ranges over a very ‘large’ space. However, it will be noted that $z_{\infty}(u;\mathfrak{z})$ only depends on $\mathfrak{z}$ through $\alpha_{\perp}^{\mathsf{T}}\overline{h}(\mathfrak{z})$: and as shown in the proof of the next result, it is always possible to find a $z_{\mathfrak{z}}\in\mathscr{M}_{\mu}$ such that $\psi(z_{\mathfrak{z}})=\alpha_{\perp}^{\mathsf{T}}\overline{h}(\mathfrak{z})$. Therefore, for the purposes of computing the full set of possible long-run effects of shocks in the model, it suffices to consider cases in which the model is in a steady state equilibrium (i.e. where $z_{\tau-1}=\dots=z_{\tau-k}=z$ for some $z\in\mathscr{M}_{\mu}$) immediately prior to the incidence of the shock.
Because $\chi^{-1}$ is not necessarily differentiable everywhere, the long-run multipliers as defined in (LABEL:multipliers) may not exist for every $z\in\mathscr{M}_{\mu}$. However, in most cases of interest, the subset $N\subset\mathscr{M}_{\mu}$ at which this occurs will be exceptionally small, corresponding e.g. in the case of piecewise affine models (see 7 below) to points exactly on the boundary between two regimes. It is also possible that the Jacobian of $\chi^{-1}$ may fail to be invertible at exceptional points: though it must be invertible almost everywhere, since $\chi$ is invertible.
@@
(i). In the linear VAR, it follows by 2.1 and D.1 that
@@
Nonetheless, let us suppose that, following Koop et al. (1996), one is also interested in ‘generalised’ impulse responses that are computed conditional on $\vec{z}_{\tau-1}=\mathfrak{z}$, but which average over all (potential) histories of the shocks subsequent to $t=\tau$. Then while we could not hope to obtain as clean a characterisation of the limiting IRF as is given above the fact that the major implication of (LABEL:multipliers), that
@@
is invariant to the state of the process suggests that this property would continue to hold for the generalised impulse responses. However, because of the possible long-range dependence of $\boldsymbol{\xi}_{t}$ on past shocks, via $\boldsymbol{\beta}_{t}$ (and hence $z_{t}$), these calculations far from straightforward, and are therefore deferred to future work.
6 Without the common row space condition
As noted in 2.2 above, a key property of models in ${\cal M}_{r}$ is that they satisfy the CRSC, the requirement that $\operatorname{im}\pi$ be an $r$-dimensional linear subspace of $\mathbb{R}^{p}$. Here we provide a simple example to substantiate the claim made above, that a violation of the CRSC will generally lead to a situation in which all the structural shocks have permanent effects, and as such the long-run identification restrictions that stem from 4.1 above are rendered unavailable.
To this end, we consider a nonlinear VAR(1), specified in VECM form as

### v2

@@ 2.2 High-level assumptions
(vi). The requirement that (LABEL:gradient) hold with $\lVert\boldsymbol{\beta}\rVert\leq\overline{b}$ implies $\boldsymbol{\theta}\circ f_{0}^{-1}$ is Lipschitz, with a Lipschitz constant that is bounded by $\overline{b}$; the converse is also true (see B.1 in B). Excepting cases where $f_{0}$ and $\boldsymbol{\theta}$ interact in a very particular way, this condition signals that our concern here is principally with models in which the equilibrium errors are defined by a function $\boldsymbol{\theta}$ for which $\boldsymbol{\theta}(z)=O(\lVert z\rVert)$ as $\lVert z\rVert\rightarrow\infty$, i.e. which exhibits at most linear asymptotic growth. For example, if $f_{0}=I_{p}$ and $\theta(z)=y-m(x)$, for $z=(y^{\mathsf{T}},x^{\mathsf{T}})^{\mathsf{T}}$, then (iii) requires $m$ to be Lipschitz, and thus that $m(x)=O(\lVert x\rVert)$ as $\lVert x\rVert\rightarrow\infty$, for $r\in\{0,\ldots,p-1\}$.
@@
For the purposes of much of this paper, the sequence of innovations $\{u_{t}\}$ may be treated as a given non-random sequence in $\mathbb{R}^{p}$, since our results are either non-asymptotic, or take limits under the assumption that the shocks cease at some finite horizon $\tau$. However, it will occasionally be useful to indicate the implications of our results in cases where the following holds.
@@ 2.3 Relationship to the linear VAR
2.4 The structural nonlinear VAR($k$)
In macroeconomics, a conventional starting point for empirical work is the structural (linear) VAR model, in which the observed series $\{z_{t}\}$ are regarded as being generated by an underlying $p$-dimensional i.i.d. sequence of structural shocks $\{\varepsilon_{t}\}$, whose elements are mutually orthogonal, and each of which have an economic interpretation (as e.g. an aggregate supply shock, a monetary policy shock, etc.). [^2] In some treatments (e.g. Lanne et al., 2017; Gouriéroux et al., 2020) the components of $\varepsilon_{t}$ are taken to be not merely uncorrelated, but mutually independent. Under non-Gaussianity, this is known to yield certain additional identifying restrictions. We emphasise that in the present work, only orthogonality between the components of $\varepsilon_{t}$ is maintained. This perspective may also be adopted within the framework of (LABEL:nlVAR), reproduced here as
$f_{0}(z_{t})=c+\sum_{i=1}^{k}f_{i}(z_{t-i})+u_{t},$ (2.1)
by supposing that there is a parametrisation $(c^{\varepsilon},\{f_{i}^{\varepsilon}\}_{i=0}^{k})$ of the model such that $\{z_{t}\}$ satisfies
$f_{0}^{\varepsilon}(z_{t})=c^{\varepsilon}+\sum_{i=1}^{k}f_{i}^{\varepsilon}(z_{t-i})+\varepsilon_{t}.$ (2.16)
We may suppose, as a convenient location and scale normalisation, that each of the elements of $\varepsilon_{t}$ have mean zero and unit variance, so that $\{\varepsilon_{t}\}$ is i.i.d. with $\mathbb{E}\varepsilon_{t}=0$ and $\mathbb{E}\varepsilon_{t}\varepsilon_{t}^{\mathsf{T}}=I_{p}$. For the purposes of this section, we shall also require $\varepsilon_{t}$ to be continuously distributed, with density $\varrho^{\varepsilon}$. (For simplicity of presentation, we also treat the order $k$ of the VAR as known, though only a known finite upper bound on the order of the VAR is required for the discussion that follows.)
The question then arises as to whether, or to what extent, the structural shocks $\varepsilon_{t}\}$ are identified by observation of $\{z_{t}\}$. To phrase this question more precisely, we first clarify how the model (LABEL:nlVAR) is parametrised. For $i\in\{0,\ldots,k\}$, let $\mathscr{F}_{i}$ denote a collection of functions $\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ to which each $f_{i}$ respectively belongs. We also specialise ERR to the case where $\{u_{t}\}$ is i.i.d., with a density $\varrho$ on $\mathbb{R}^{p}$ that belongs to some set $\mathscr{R}$, normalised to have $\mathbb{E}u_{t}=0$ and $\mathbb{E}u_{t}u_{t}^{\mathsf{T}}=I_{p}$. The parameter space for (LABEL:nlVAR) is then
$\mathscr{P}\coloneqq\{(\tilde{c},\{\tilde{f}_{i}\}_{i=0}^{k},\tilde{\varrho})\mid\tilde{c}\in\mathbb{R},\ \tilde{f}_{i}\in\mathscr{F}_{i}\text{ for }i\in\{0,\ldots,k\},\ \tilde{\varrho}\in\mathscr{R}\}.$
Under the regularity conditions given in A, each $(\tilde{c},\{\tilde{f}_{i}\}_{i=0}^{k},\tilde{\varrho})\in\mathscr{P}$ completely specifies the density of $z_{t}$ conditional on $\vec{z}_{t-1}=(z_{t-1}^{\mathsf{T}},\ldots,z_{t-k}^{\mathsf{T}})^{\mathsf{T}}$. Following Matzkin (2008, pp. 949f.), we say that two points in $\mathscr{P}$ are observationally equivalent if they imply exactly the same conditional density function, and thus the same likelihood for $\{z_{t}\}_{t=1}^{n}$, conditional on the initial values $\vec{z}_{0}$ (which as per CVAR we take to be arbitrary, and thus uninformative about the model parameters).
Let $\Upsilon\in\mathbb{R}^{p\times p}$ be an orthogonal matrix, and suppose that $\mathscr{R}$ is such that the density $\tilde{\varrho}$ of $\tilde{u_{t}}\coloneqq\Upsilon\varepsilon_{t}$ also lies in $\mathscr{R}$. Then it is evident from pre-multiplying (LABEL:struct-form) through by $\Upsilon$, to obtain
$\tilde{f}_{0}(z_{t})\coloneqq\Upsilon f_{0}^{\varepsilon}(z_{t})=\Upsilon c^{\varepsilon}+\sum_{i=1}^{k}\Upsilon f_{i}^{\varepsilon}(z_{t-i})+\Upsilon\varepsilon_{t}\eqqcolon\tilde{c}+\sum_{i=1}^{k}\tilde{f}_{i}(z_{t-i})+\tilde{u}_{t},$
that $(\tilde{c},\{\tilde{f}_{i}\}_{i=0}^{k},\tilde{\varrho})$ must be observationally equivalent to $(c^{\varepsilon},\{f_{i}^{\varepsilon}\}_{i=0}^{k},\varrho^{\varepsilon})$. Thus, as is familiar from the linear structural VAR, the model parameters and the structural shocks are at best identified up to an unknown orthogonal transformation.
Remarkably, despite the much broader class of nonlinear transformations permitted by (LABEL:nlVAR), such orthogonal transformations delineate the only loci of non-identification in the nonlinear VAR. By A.1 in A, under certain conditions on the sets $\{\mathscr{F}_{i}\}_{i=0}^{k}$ and $\mathscr{R}$, and the parameters generating $\{z_{t}\}$, there exists a $\varrho\in\mathscr{R}$ such that the parameters $(c,\{f_{i}\}_{i=0}^{k},\varrho)$ of (LABEL:nlVAR) are observationally equivalent to those $(c^{\varepsilon},\{f_{i}^{\varepsilon}\}_{i=0}^{k},\varrho^{\varepsilon})$ of (LABEL:struct-form), if and only if there is an orthogonal matrix $\Upsilon\in\mathbb{R}^{p\times p}$ such that
$c=\Upsilon c^{\varepsilon}\text{ and }f_{i}(z)=\Upsilon f_{i}^{\varepsilon}(z),\ \forall z\in\mathbb{R}^{p},\ i\in\{0,\ldots,k\},$ (2.17)
and thus the shocks $\{u_{t}\}$ in (LABEL:nlVAR) satisfy
$u_{t}=f_{0}(z_{t})-c-\sum_{i=1}^{k}f_{i}(z_{t-i})=\Upsilon\left[f_{0}^{\varepsilon}(z_{t})-c^{\varepsilon}-\sum_{i=1}^{k}f_{i}^{\varepsilon}(z_{t-i})\right]=\Upsilon\varepsilon_{t}.$ (2.18)
(For a discussion of the conditions under which the preceding holds, see A: these are essentially weak smoothness conditions on the elements of $\{\mathscr{F}_{i}\}_{i=0}^{k}$ and $\mathscr{R}$, together with the requirement that $\{f_{i}^{\varepsilon}\}_{i=1}^{k}$be such as to ensure sufficient dependence of the r.h.s. of the model on lags of $z_{t}$.) Conversely, should (LABEL:identified-set) fail to hold, then there will be at least some realisations of $\{z_{t}\}$ for which the likelihoods of $(c,\{f_{i}\}_{i=0}^{k},\varrho)$ and $(c^{\varepsilon},\{f_{i}^{\varepsilon}\}_{i=0}^{k},\varrho^{\varepsilon})$ will be distinct, and so the data is to this extent informative about these two alternative parametrisations of the model. [^3] We do not claim, on the basis of A.1, that the parameters of the model are consistently estimable up to an orthogonal transformation. While it seems reasonable to suppose that consistent nonparametric estimation of the model would be possible (under regularity conditions) if $\{z_{t}\}$ is stationary and ergodic, the usual connection between identification and consistent estimation is attenuated when $\{z_{t}\}$ possesses some stochastic (or indeed, deterministic) trends, because of the non-recurrence of those trends in higher dimensions (see Bingham, 2001, Sec. 6; Gao and Phillips, 2013, p. 62). Consistent estimation of the model parameters (up to $\Upsilon$) would in such cases likely require further restrictions on the functional form of $f_{i}$.
Accordingly, when subsequently discussing the problem of structural shock identification, we shall suppose that the shocks $u_{t}$ appearing in (LABEL:nlVAR) are related to $\varepsilon_{t}$ in precisely the manner of (LABEL:u-to-epsilon), yielding the structural nonlinear VAR model
where $\Upsilon\in\mathbb{R}^{p\times p}$ is an orthogonal matrix. In other words, on the strength of A.1, we shall regard observation of $\{z_{t}\}$ as being sufficient to identify the model parameters up to (and only up to) $\Upsilon$, and thence the structural shocks up to transformation by this same matrix. To identify the structural shocks, we thus seek additional restrictions sufficient to pin down (at least some of) the unknown elements of $\Upsilon$, such as will be provided by the long-run identifying restrictions developed in 4.2 below.
3 Granger–Johansen representation theory
@@
Before providing our extension of the GJRT to the general setting of (LABEL:nlVAR), we first note that the nonlinear counterpart of $\boldsymbol{\xi}_{t}$ will continue to follow an autoregression of the form (LABEL:eqerr-linear), but now with time-varying coefficients, in particular with $\boldsymbol{\beta}=\boldsymbol{\beta}_{t}$ now depending on the level of $z_{t}$. While such processes cannot be stationary, we still have a well-defined notion of stability for such processes, which is sufficient to ensure that $\boldsymbol{\xi}_{t}$ remains of strictly smaller stochastic order than the common trends.
@@
We now state our counterpart of the Granger–Johansen representation theorem for the model (LABEL:nlVAR), which constitutes the main result of this paper.
@@
$S_{\chi}^{\mathsf{T}}\coloneqq\begin{bmatrix}-\alpha_{\perp}^{\mathsf{T}}&0\\ 0&I_{r}\end{bmatrix}\begin{bmatrix}0_{p\times r}&I_{p}&\cdots&I_{p}\\ I_{r}&0_{r\times p}&\cdots&0_{r\times p}\end{bmatrix},$
@@
by E.1 in E, which agrees precisely with Johansen (1995, Ch. 4).
@@
4 Consequences of the representation theorem
4.1 Attractor spaces
@@
where the fact that $\mathscr{M}_{\mu}$ is indeed a $q$-dimensional manifold follows from the final equality, since $\chi:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ is a homeomorphism.
@@
${\cal S}\subset\mathbb{R}^{p}$ is stable if for every $(u,\mathfrak{z})\in\mathbb{R}^{p}\times\mathbb{R}^{kp}$, there exists a $z_{\infty}\in{\cal S}$ such that $z_{t}(u,\mathfrak{z})\rightarrow z_{\infty}\in{\cal S}$. $z_{\infty}\in\mathbb{R}^{p}$ is a non-trivial attractor if for every $\mathfrak{z}\in\mathbb{R}^{kp}$, there exists a non-empty $\mathscr{U}(z_{\infty};\mathfrak{z})\subset\mathbb{R}^{p}$ such that $z_{t}(u,\mathfrak{z})\rightarrow z_{\infty}$ for all $u\in\mathscr{U}(z_{\infty};\mathfrak{z})$; we term $\mathscr{U}(z_{\infty};\mathfrak{z})$ the domain of attraction for $z_{\infty}$, given $\mathfrak{z}$. ${\cal S}$ is strictly stable if it is stable and contains only non-trivial attractors.
In other words, if $\mathscr{M}_{\mu}$ is strictly stable, then whatever the current state $\vec{z}_{\tau-1}=\mathfrak{z}$ of the model and the given limiting $z_{\infty}\in\mathscr{M}_{\mu}$, there is always a value for the $\tau$-dated shock $u_{\tau}$ that would lead $z_{t}$ to converge to $z_{\infty}$; every element in $\mathscr{M}_{\mu}$ may thus ultimately be ‘reached’ from $\vec{z}_{\tau-1}$. (Note that this is not so trivial a matter as choosing $u_{\tau}=u$ such that $z_{\tau}=z_{\infty}$ immediately, since what is required is that $z_{t}$ converge to a steady state equilibrium at $z_{\infty}$.) Since $\{\boldsymbol{\xi}_{t}\}$ is exponentially stable by 3.1, it follows that $\boldsymbol{\xi}_{t}\rightarrow 0$ as $t\rightarrow\infty$ (see the discussion following 3.1 above). Hence, noting that $\sum_{s=\tau}^{t}u_{s}=u_{\tau}=u$, we have by that result that
@@
Because the r.h.s. depends on $\mathfrak{z}$ only through $\overline{h}(\mathfrak{z})$, and as $u\in\mathbb{R}^{p}$ varies $\alpha_{\perp}^{\mathsf{T}}[\overline{h}(\mathfrak{z})+u]$ ranges freely over $\mathbb{R}^{q}$, we can induce $z_{\infty}(u;\mathfrak{z})$ to take any desired value in $\mathscr{M}_{\mu}$. We thus obtain the following.
@@
4.2 Long-run identifying restrictions
As a direct consequence of the preceding, we obtain the following characterisation of the linear combinations of the shocks $u_{t}$ in (LABEL:nlVAR) that do not have permanent effects.
@@ 4.2 Long-run identifying restrictions
The significance of this result may be explained as follows. Suppose that, as in (LABEL:nlSVAR) above, we have the nonlinear structural VAR
belonging to class ${\cal M}_{r}$. As noted in 2.4, the parameters of this model are identified up to the unknown orthogonal matrix $\Upsilon\in\mathbb{R}^{p\times p}$, about which the data is entirely uninformative, and thus we seek additional restrictions that would pin down (at least some of) the elements of $\Upsilon$. To that end, suppose we partition the structural shocks as
@@
where $\varepsilon_{(1),t}$ takes values in $\mathbb{R}^{m}$, and collects the $m\in\{1,\ldots,r\}$ structural shocks that are regarded as having no permanent effect on $z_{t}$. Then 4.1 implies that
$\Upsilon\begin{bmatrix}I_{m}\\ 0_{(p-m)\times m}\end{bmatrix}\subset\operatorname{sp}\alpha\implies\alpha_{\perp}^{\mathsf{T}}\Upsilon\begin{bmatrix}I_{m}\\ 0_{(p-m)\times m}\end{bmatrix}=0,$ (4.5)
which provides $qm$ ‘long-run identifying restrictions’ on $\Upsilon$ [^4] Note that 4.1, or equivalently (LABEL:lr-ident), implies that at most $r$ elements of $\varepsilon_{t}$ may be such that the corresponding column of $\Upsilon$ lies in $\operatorname{sp}\alpha$, i.e. up to $r$ structural shocks may have purely transitory effects. As is familiar from the linear structural VAR, there is nothing here to preclude e.g. all shocks from having permanent effects; the number (and identities) of the $m\in\{0,\ldots,r\}$ structural shocks having purely transitory effects will thus depend on the identifying conditions that define those shocks.
As discussed further in 5 below, the availability of these long-run identifying restrictions is largely a consequence of the common row space condition, and provides one of the principal motivations for maintaining that condition in our model. Remarkably, despite the nonlinearity permitted by (LABEL:nlSVAR), the identifying restrictions (LABEL:lr-ident) take exactly the same form as in a linear SVAR (see Kilian and Lütkepohl, 2017, Sec. 10.2).
4.3 Long-run multipliers
Having obtained a characterisation of the attractor space $\mathscr{M}_{\mu}$ for $\{z_{t}\}$, we may also say something more about the permanent effect of a shock at time $t=\tau$. The limiting impulse responses or long-run multipliers can be computed by differentiating $z_{\infty}(u;\mathfrak{z})$ with respect to $u$, i.e. by computing the Jacobian
@@ 4.3 Long-run multipliers
In view of (LABEL:nlSVAR), postmultiplying this by $\Upsilon$ yields the long-run multipliers with respect to the structural shocks $\varepsilon_{t}$, at time $t=\tau$.
One concern we might have here, in the general nonlinear case, is the apparent dependence of these long-run multipliers on $\vec{z}_{t-1}=\mathfrak{z}\in\mathbb{R}^{kp}$, which potentially ranges over a very ‘large’ ($kp$-dimensional) space. However, it will be noted that $z_{\infty}(u;\mathfrak{z})$ only depends on $\mathfrak{z}$ through $\alpha_{\perp}^{\mathsf{T}}\overline{h}(\mathfrak{z})$: and as shown in the proof of the next result, it is always possible to find a $z_{\mathfrak{z}}\in\mathscr{M}_{\mu}$ such that $\psi(z_{\mathfrak{z}})=\alpha_{\perp}^{\mathsf{T}}\overline{h}(\mathfrak{z})$. Therefore, for the purposes of computing the full set of possible long-run effects of shocks in the model, it suffices to consider cases in which the model is in a steady state equilibrium (i.e. where $z_{\tau-1}=\dots=z_{\tau-k}=z$ for some $z\in\mathscr{M}_{\mu}$) immediately prior to the incidence of the shock.
Because $\chi^{-1}$ is not necessarily differentiable everywhere, the long-run multipliers as defined in (LABEL:multipliers) may not exist for every $z\in\mathscr{M}_{\mu}$. However, in most cases of interest, the subset $N\subset\mathscr{M}_{\mu}$ at which this occurs will be exceptionally small, corresponding e.g. in the case of piecewise affine models (see 6 below) to points exactly on the boundary between two regimes. It is also possible that the Jacobian of $\chi^{-1}$ may fail to be invertible at exceptional points: though it must be invertible almost everywhere, since $\chi$ is invertible.
@@
(i). In the linear VAR, it follows by 2.1 and E.1 that
@@
Nonetheless, let us suppose that, following Koop et al. (1996), one is also interested in ‘generalised’ impulse responses that are computed conditional on $\vec{z}_{\tau-1}=\mathfrak{z}$, but which average over all (potential) histories of the shocks subsequent to $t=\tau$. Then while we could not hope to obtain as clean a characterisation of the limiting IRF as is given by the preceding theorem, the fact that the major implication of (LABEL:multipliers), that
@@
is invariant to the state of the process suggests that this property would continue to hold for the generalised impulse responses. However, because of the possible long-range dependence of $\boldsymbol{\xi}_{t}$ on past shocks, via $\boldsymbol{\beta}_{t}$ (and hence $z_{t}$), these calculations are far from straightforward, and are therefore deferred to future work.
5 The common row space condition
As introduced in 2.1 above, a key characteristic of models in class ${\cal M}_{r}$ is the common row space condition (CRSC), that $\operatorname{im}\pi=\{\pi(z)\mid z\in\mathbb{R}^{p}\}$ is an $r$-dimensional linear subspace of $\mathbb{R}^{p}$; we noted there that no such restriction is imposed on $\ker\pi$. By contrast, the existing literature on nonlinear VECM models effectively reverses this condition by requiring that $\ker\pi$ be a $q$-dimensional subspace of $\mathbb{R}^{p}$, without necessarily restricting $\operatorname{im}\pi$ (see in particular Kristensen and Rahbek, 2010). Loosely speaking, if we suppose that $\pi$ may be decomposed, for each $z\in\mathbb{R}^{p}$, as
$\pi(z)=\alpha(z)\beta(z)^{\mathsf{T}}$ (5.1)
where $\alpha,\beta:\mathbb{R}^{p}\rightarrow\mathbb{R}^{r}$, then the CRSC permits taking $\alpha(z)=\alpha$, if $\beta(z)$ is appropriately normalised, but leaves $\beta(z)$ otherwise unrestricted; [^5] If $\beta(z)$ is instead normalised in some other way, such as e.g. $\beta(z)^{\mathsf{T}}=[I_{r},-A(z)]$, then $\alpha(z)$ may indeed vary with $z$. In this sense, it is not quite correct to regard the CRSC as requiring the loadings on the equilibrium errors to be constant, but rather that those loadings should lie in a fixed $r$-dimensional linear subspace; there may be nonlinear adjustment towards equilibrium, provided that it occurs along this subspace. whereas in the nonlinear VECM literature, $\beta(z)=\beta$ is constant, with the result that there is a globally linear cointegrating space. The representation theory developed in the nonlinear VECM literature is thus strictly complementary to the present work. (Note that even the case where $\beta(z)=\beta$ in our framework is not encompassed by that literature, because we allow the dynamics of the implied VECM (LABEL:VECM-first) to depend on the level of $z_{t}$, whereas that literature requires that these be governed entirely by the stationary equilibrium errors $\beta^{\mathsf{T}}z_{t}$ and differences $\Delta z_{t}$.)
Recall that in the linear structural VAR model, common stochastic trends arise because some subset of the structural shocks (say, $q$ in number) have permanent effects on $z_{t}$, while the remaining $r=p-q$ shocks have only transitory effects. This provides not only an underpinning economic explanation for the patterns of long-run co-movement present in the data, but also a fruitful source of long-run identifying restrictions (following Blanchard and Quah, 1989, and King et al., 1991). The results developed in Sections 3 and 4 above show that these properties are preserved when we generalise the cointegrated linear SVAR to a nonlinear SVAR (LABEL:nlSVAR) belonging to class ${\cal M}_{r}$. However, as the example developed in this section illustrates, these properties are highly sensitive to departures from the CRSC, even when the assumption of a globally linear cointegrating space is maintained (i.e. when $\beta(z)=\beta$ for all $z\in\mathbb{R}^{p}$ in (LABEL:factor) above). In models where the CRSC fails, it will generally be the case that the direction in which an impulse $u_{t}=\delta$ has no permanent effect on $z_{t}$ will vary with the magnitude $\delta$. Thus, if we want to work with nonlinear SVAR models in which structural shocks may be distinguished on the basis of their long-run effects – and in which, as a corollary, common stochastic trends arise because only a subset of these shocks have permanent effects – then it would appear that the CRSC cannot be easily dispensed with.
To illustrate the consequences of a failure of the CRSC, we consider a nonlinear VAR(1), specified in VECM form as